Session: Feb 2005 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (scalambra at Fri Mar 4 09:02:35 2005) Failure : RECO FS
Mark5 recording problem. Error message was recording throttle. This caused the field system to hang. Anly first 70 scans were recorded.

Jodrell_Bank (Alastair Gunn at Mon Mar 7 12:11:25 2005) Minor failures : POIN
The Lovell Telescope was used. The Lovell telescope was used. Drive tripped between 20:21 and 20:22 UT. No other known problems.

Cambridge (Alastair Gunn at Mon Mar 7 12:12:51 2005) Success :
No known problems.

Westerbork (Tony Foley at Mon Mar 7 15:14:10 2005) Minor failures : WIND HIGH
Wb used in phased array mode. Using a 'tied array'with only one telescope (RT7). Still lots of snow in the dishes, but we tried to clean some snow from RT7.

Noto (P. Cassaro at Mon Mar 7 15:33:09 2005) Success :
No known problems

Effelsberg (A. Kraus at Fri Mar 11 12:34:47 2005) Minor failures : WIND
stopped due to snowfall: UT9515-0640, UT0710-0803.